Clare wants to buy a used wide-screen TV that sells for $2,000 cash. She will make a down payment of $1,000 and then six monthly payments of $200. What is the full installment price and the annual interest rate?

Answer by rfer(16322)

2000-1000=1000
200*6=1200
1000+1200=$2200 total
----------
(200/1000)/2=0.10=10%